OpenStudy (anonymous):

Find the missing parts of the triangle. (Find angles to the nearest hundredth of a degree.) a = 6.1 in. b = 13.6 in. c = 16.0 in. Answer a A = 21.92°, B = 56.34°, C = 101.74° b A = 19.92°, B = 56.34°, C = 103.74° c A = 23.92°, B = 54.34°, C = 101.74° d No triangle satisfies the given conditions.

OpenStudy (anonymous):

well wait.. cant you use the answers for your own good?

OpenStudy (alexwee123):

this time USE LAW OF COSINES

OpenStudy (anonymous):

wait nvm. it doesnt tell u an angle haha

OpenStudy (anonymous):

\[Cos \theta(C) =\frac{a^{2}+b^{2}-c^{2}}{2ab}\]

OpenStudy (alexwee123):

there were so many trig questions today :o

OpenStudy (anonymous):

for angle B \[\cos \theta (B) =\frac{ a^{2}+c^{2}-b^{2}}{2ac}\]

OpenStudy (anonymous):

A is the answer, btw @helloimjared u have to work this out

OpenStudy (anonymous):

You saying u want me to draw it on here and get the answer?